# All Patches Matter, More Patches Better: Enhance AI-Generated Image Detection via Panoptic Patch Learning

### v1

@@ 5.1 Comparison with other methods
Tab. 2 reports the comparison on DRCT. We obtain the results from DRCT [1].The results indicate the following: (1) PPL consistently achieves the highest average accuracy with the lowest std. (2) DRCT shows poor detection performance on the SDXL-related subset, while PPL demonstrates a more balanced performance across various subsets. Overall comparisons in Setting-I indicate that PPL has a greater generalizing ability across different generation models.
@@
Fig. 8 compares the performance of PPL with other methods in the Chameleon dataset, utilizing the entire GenImage training set. Performance metrics for all the methods listed are cited from the paper Yan et al. [36]. The results indicate that most existing methods struggle to achieve an accuracy of approximately 55%, which only marginally exceeds the accuracy of random guessing (50%). In contrast, our method achieved an accuracy of 70% on Chameleon, surpassing the state-of-the-art (SOTA) method by 5%. This demonstrates a superior generalization ability when faced with higher-quality and fine-tuned versions of generative models.
5.2 Comparisons on Robustness
We conduct a series of robustness experiments on GenImage to verify the reliability of our method against image corruption. Additionally, we perform robustness experiments against random masking, demonstrating PPL’s ability to leverage a greater number of available patches. For a fair comparison, we reimplemented UnivFD [19] using the same basic data augmentation (cropping, rotation, JPEG compression, etc.) as those employed in our approach.
@@ 5.2 Comparisons on Robustness
We assess the accuracy of our method under JPEG compression(quality factor = 100, 90, 80, 70, 60) and Gaussian blur (deviation degree = 0.6, 0.8, 1.0, 1.2, 1.4) on the SDv1.4 subset of GenImage dataset. Fig. 9 illustrates that both backbones of our approach sustain high accuracy even under extreme JPEG compression and Gaussian blur, maintaining an accuracy of approximately 90%.
@@
We assess the recall rate of methods by gradually masking out portions of the input images to evaluate the models’ ability to detect fake artifacts from a greater number of patches. Due to the varying initial recall values among different methods, we measure the decline in recall as the ratio of the decrease in recall to the original value. Fig. 10 demonstrates that PPL exhibits the best robustness against masking, with only a 5% decline under extreme masking conditions, confirming that PPL reduces over-reliance on specific patches.
5.3 Ablation Studies
To investigate the impact of each component and hyperparameter in PPL, we conducted a series of ablation studies. Unless otherwise specified, we used CLIP as the backbone model, and trained on the SDv1.4 subset of GenImage. We then compared the overall accuracy (mACC) of each result obtained on the whole GenImage dataset.
Ablation on the impact of each module.
| Lora | RPR | PCL | mAcc |
@@ 5.3 Ablation Studies
Ablation on hyperparameters.
Tab. 4 illustrates the impact of key hyper parameters in our framework. Based on the experimental results, we draw the following conclusions: (1) The weight setting of contrastive loss is highly sensitive, with an optimal value of 0.3. (2) Model performance improves with the increase in mixing data seen during training, suggesting the effectiveness of mixing data. (3) General accuracy improves as the mixing ratio increases from 10% to 50%, peaking at 50%, but experiences a significant decline at 90%. This suggests that replacing too many patches in an AI-generated image with real ones sets an overly challenging goal for the model, which hurts its performance.
| mAcc | |
| 0.1 | 90.9 |
| 0.3 | 97.2 |
| 0.5 | 95.8 |
| 0.7 | 95.1 |
| 0.9 | 91.5 |
| mAcc | |
| 10 | 91.7 |
| 30 | 94.7 |
| 50 | 97.2 |
| 70 | 94.4 |
| 90 | 77.4 |
| mAcc | |
| 2.5 | 94.9 |
| 7.5 | 95.9 |
| 12.5 | 96.1 |
| 17.5 | 95.4 |
| 22.5 | 97.2 |
@@ 6 Conclusion
Our work begins with a discussion on the nature of the AIGI detection problem, which can be concluded as, ‘All Patches Matter, More Patches Better.’ However, our observations indicate that existing detectors are unable to fully take advantage of all patches in an AI-generated image. To address this issue, we propose a Random Patch Replacement augmentation combined with a Patch-wise Contrastive Learning strategy. This approach effectively prevents the model from becoming a lazy learner and enhances the utilization of every patch. We achieve state-of-the-art performance on several well-known academic datasets across two settings: one restricts the training set to evaluate generalization ability, while the other includes more challenging test cases without limiting the model’s training set. Although we have made significant progress, there remains room for improvement in the second setting; thus, future work may focus on developing higher-quality training sets to enhance performance on contemporary real-world benchmarks.

### v2

@@ 5.1 Comparison with other methods
Tab. 2 reports the comparison on DRCT. We obtain the results from DRCT [1].The results indicate the following: (1) PPL consistently achieves the highest average accuracy with the lowest std. (2) DRCT shows poor detection performance on SDXL-related subsets, while PPL demonstrates a more balanced performance across various subsets. Overall comparisons in Setting-I indicate that PPL has a greater generalizing ability across different generation models.
@@
Fig. 3 compares the performance of PPL with other methods in the Chameleon dataset, utilizing the entire GenImage training set. Performance metrics for all the methods listed are cited from the paper [36]. The results indicate that most existing methods struggle to achieve an accuracy of approximately 55%, which only marginally exceeds the accuracy of random guessing (50%). In contrast, our method achieved an accuracy of 70% on Chameleon, surpassing the SOTA method by 5%. This demonstrates a superior generalization ability when faced with higher-quality and fine-tuned versions of generative models.
5.2 Robustness Studies
We conduct a series of robustness experiments on GenImage to verify the reliability of our method against image corruption. Additionally, we perform robustness experiments against random masking, demonstrating PPL’s ability to leverage a greater number of available patches. For a fair comparison, we re-implemented UnivFD [20] using the same basic data augmentation (cropping, rotation, JPEG compression, etc.) as those employed in our approach.
@@ 5.2 Robustness Studies
We assess the accuracy of our method under JPEG compression(quality factor = 100, 90, 80, 70, 60), Gaussian blur (standard deviation = 0.5, 1.0, 1.5, 2.0, 2.5) and resizing (scale = 0.6, 0.8, 1.0, 1.2, 1.4) on the SDv1.4 subset of GenImage dataset. Fig. 6 illustrates that both backbones of our approach sustain high accuracy even under extreme JPEG compression and Gaussian blur, maintaining an accuracy of approximately 90%.
@@
We evaluate the recall rates of various methods by progressively masking portions of the input images to assess each model’s ability to detect synthetic artifacts by leveraging a greater number of patches. To account for varying initial recall values among methods, we quantify recall decline as the ratio of the decrease in recall to the original value. As shown in Fig. 6, PPL demonstrates the highest robustness against masking, with only a 5% decline under extreme masking conditions. This result confirms that PPL reduces over-reliance on specific patches.
Robustness studies demonstrate that reducing reliance on a few local patches and paying attention to patches from entire image not only improves performance metrics on academic datasets, but also enhances robustness against perturbations.
5.3 Ablation Study
To investigate the impact of each component in PPL, we conducted a series of ablation studies. Unless otherwise specified, we used CLIP as the backbone model, and trained on the SDv1.4 subset of GenImage. We then compared the overall accuracy (mACC) of each result obtained on the whole GenImage dataset.
| Lora | RPR | PCL | mAcc (%) |
@@ 5.3 Ablation Study
Ablation on the impact of each module.
6 Limitations and Future Work
In this work, we note that all patches of a fully AI-generated image may contain useful synthetic artifacts and propose a patch-wise learning strategy to leverage information from all patches. Using more advanced methods and larger models to identify synthetic patches would be of our future interest and could further improve detection performance. Furthermore, with the advancement of T2I models and the growing AIGC research community, state-of-the-art models now surpass the quality of images in existing academic benchmarks. Future works would aim to build datasets close to real-world scenarios and find solutions to them, enhancing the practical utility of AIGI detection.
@@ 7 Conclusion
Our work is based on the nature of the AIGI detection problem, which can be concluded as "All Patches Matter, More Patches Better." However, our observations indicate that existing detectors are unable to fully take advantage of all patches in an AI-generated image. To address this issue, we propose a Random Patch Replacement augmentation combined with a Patch-wise Contrastive Learning strategy. This approach effectively prevents the model from becoming a lazy learner and enhances the utilization of every patch. We achieve state-of-the-art performance on several well-known academic datasets across two settings: one restricts the training set to evaluate generalization ability, while the other includes more challenging test cases without limiting the model’s training set. The outstanding performance achieved in both settings supports our findings and proves the efficacy of the proposed learning framework.